hi , thank you for your time. I'm in grade 11 math and my name is Rita.
My question is:
Resolve algalberqy the systeme of equation f(x)=3x+2 and f(x)=2x-3
and also
thank you very much

Answer by richwmiller(17219)

What is "algalberqy" ?
Some website rules you won't like:
You are limited to one problem per submission. One problem at a time
No more than 4 submissions daily 4 problems a day
No similar problems. Different type problems only
f(x)=3x+2 and f(x)=2x-3
set them equal to each other
3x+2=2x-3
x=5
so plug in 5 for x in either equation
2*5-3=y
7=y
(5,7)
done